# cell-space-top-limit = 2pt, cell-space-bottom-limit = 1.5pt Giant orbital magnetization in two-dimensional materials

### v1

@@ Abstract
Orbital magnetization typically plays a minor role in compounds where the magnetic properties are governed by transition metal elements. However, in some cases, the orbital magnetization may be fully unquenched, which can have dramatic consequences for magnetic anisotropy and various magnetic response properties. In the present work, we start by summarizing how unquenched orbital moments arise from particular combinations of crystal field splitting and orbital filling. We exemplify this for the cases of two-dimensional (2D) VI3 and FePS3, and show that Hubbard corrections as well as self-consistent spin-orbit coupling are crucial ingredients for predicting correct orbital moments from first principles calculations. We then search the Computational 2D Materials Database (C2DB) for monolayers having tetrahedral or octahedral crystal field splitting of transition metal -states and orbital occupancy that is expected to lead to large orbital moments. We identify 112 monolayers with octahedral crystal field splitting and 62 monolayers with tetrahedral crystal field splitting and for materials with partially filled bands, we verify that inclusion of Hubbard corrections as well as self-consistent spin-orbit coupling typically increases the magnitude of predicted orbital moments by an order of magnitude.
@@ I Introduction
Magnetism at the atomic level can be understood through contributions from spin and orbital magnetic moments. Spin magnetization, like electric charge, is an intrinsic property of the electron and can be described in terms of a local magnetization density . This is a well-defined field that characterizes the many-body ground state and may, for example, be calculated in the framework of density functional theory (DFT) [1]. Moreover, the response of the spin magnetization under a dynamic external field is governed by the fundamental (spin) magnetic excitations, which can be calculated from either time-dependent DFT [2, 3, 4, 5, 6, 7, 8, 9] or many-body perturbation theory [10, 11, 12]. The magnetization density is typically localized in the vicinity of atoms and the dynamics may often be accurately approximated by a simple site-based model with inter-site interactions calculated from first principles. Such models are routinely used to predict magnon energies as well as critical temperatures for magnetic order and provide a rather direct link to inelastic neutron scattering experiments, which are typically interpreted in terms of site-based models.
@@ III Computational details
Each material was subjected to two different ground state calculations using the open-source electronic structure package GPAW [45, 46, 47] using the projector augmented-wave method [48]. The first calculation was performed with spin-polarized collinear DFT using LSDA and SOC added non-self-consistently subsequently. The second calculation used non-collinear DFT with LSDA, a Hubbard correction on the transition metal -orbitals with eV and SOC included self-consistently. For each calculation, the local orbital moments were calculated with the projector augmented-wave atom-centered approximation (PAW-ACA) [40]. All calculations were performed using a plane wave basis with a cutoff energy of 800 eV and a -point density of at least 12 Å.

### v2

# Giant orbital magnetization in two-dimensional materials

@@ Abstract
Orbital magnetization typically plays a minor role in compounds where the magnetic properties are governed by transition metal elements. However, in some cases, the orbital magnetization may be fully unquenched, which can have dramatic consequences for magnetic anisotropy and various magnetic response properties. In the present work, we start by summarizing how unquenched orbital moments arise from particular combinations of crystal field splitting and orbital filling. We exemplify this for the cases of two-dimensional (2D) VI3 and FePS3, and show that Hubbard corrections as well as self-consistent spin-orbit coupling are crucial ingredients for predicting correct orbital moments from first principles calculations. We then search the Computational 2D Materials Database (C2DB) for monolayers having tetrahedral or octahedral crystal field splitting of transition metal -states and orbital occupancy that is expected to lead to large orbital moments. We identify 112 monolayers with octahedral crystal field splitting and 62 monolayers with tetrahedral crystal field splitting and for materials with partially-filled bands, we verify that inclusion of Hubbard corrections as well as self-consistent spin-orbit coupling typically increases the magnitude of predicted orbital moments by an order of magnitude.
@@ I Introduction
Magnetism at the atomic level can be understood through contributions from spin and orbital magnetic moments. Spin magnetization, like electric charge, is an intrinsic property of the electron and can be described in terms of a local magnetization density . This is a well-defined field that characterizes the many-body ground state and may, for example, be calculated in the framework of density functional theory (DFT) [1]. Moreover, the response of the spin magnetization under a dynamic external field is governed by the fundamental (spin) magnetic excitations, which can be calculated from either many-body perturbation theory [2, 3, 4] or time-dependent DFT as has been applied extensively on simple bulk magnets such as Fe, Co and Ni [5, 6, 7, 8, 9, 10] or more complicated crystals such as bulk NiO and MnO [11] and monolayer Fe3GeTe2 [12]. The magnetization density is typically localized in the vicinity of atoms and the dynamics may often be accurately approximated by a simple site-based model with inter-site interactions calculated from first principles. Such models are routinely used to predict magnon energies as well as critical temperatures for magnetic order and provide a rather direct link to inelastic neutron scattering experiments, which are typically interpreted in terms of site-based models.
@@ II.1 Crystal field splitting and the filling of orbitals
The physical differences between eigenstates of the angular momentum operator and states like can be further highlighted by considering their respective loop-currents as represented by the winding of the complex phase. As Fig. 2 shows, eigenstates of an angular momentum operator such as of the operator represent looping currents that follow perfectly circular paths which lie in planes perpendicular to the -axis whereas the looping current of must follow a more undulating path, maximizing the distance between the electron and the repulsive electron clouds near the ligands as the states do by construction.
@@ III Computational details
Each material was subjected to two different ground state calculations using the open-source electronic structure package GPAW [45, 46, 47] using the projector augmented-wave method [48]. The first calculation was performed with spin-polarized collinear DFT using LSDA and SOC added non-self-consistently subsequently. The second calculation used non-collinear DFT with LSDA, a Hubbard correction on the transition metal -orbitals with eV and SOC included self-consistently. Note that the choice of eV is largely arbitrary as the main purpose of the Hubbard penalty is to be strong enough to split the partially-filled states. Nevertheless, it has been demonstrated for FePS3 and VI3 that increasing the Hubbard penalty beyond eV does not alter the atomic orbital moments significantly [40]. Furthermore, the initial spin configuration for each non-collinear DFT calculation was chosen as a collinear ferromagnet pointing along the previously determined easy axis of each crystal [49]. Reasonably, one might suspect that ferromagnetic and anti-ferromagnetic states may yield different magnitudes of the orbital moments; however, for the prototypical FePS3 crystal this was not the case, and since giant orbital moments arise from the local ligand environment, the long-range magnetic order should not play a major role on the magnitude of the orbital moments. For both collinear and non-collinear DFT calculation, the local orbital moments were calculated with the projector augmented-wave atom-centered approximation (PAW-ACA) [40]. All calculations were performed using a plane wave basis with a cutoff energy of 800 eV and a -point density of at least 12 Å.
@@ IV Results
Since these unquenched orbital moments arise from the ligand geometry around a transition metal and since the SOC couple the directions of spin and orbital magnetic moments, the presence of unquenched orbital moments may signify an unusually large magnetic anisotropy as the orbital moment firmly pins the direction of spin moment to the lattice. This enhancement of the magnetic anisotropy shall be shown directly through constrained-spin DFT of the two-dimensional FePS3 crystal. Through constrained-spin DFT, the direction of spins near the Fe atoms can be forced to point along directions varying between (pointing out of the crystal plane) and (pointing along the crystal plane). The resulting increase in the potential energy of the crystal as the spin direction is forced further in-plane is displayed in Fig. 5. Additionally shown is also the components of the orbital magnetic moment at the Fe atoms which continue to point predominantly out-of-plane even as the direction of the local spin moments are completely in-plane. This increase in the crystal energy coupled with the mostly static direction and magnitude of the local orbital magnetic moments exemplify the connection between magnetic anisotropy and giant orbital magnetic moments.